the legs of a ladder are of equal length and are 1 meter apart. the slant height of the ladder is 2.4 metres. how high above the ground does the ladder reach?
Answer by stanbon(75887)

The legs of a ladder are of equal length and are 1 meter apart. the slant height of the ladder is 2.4 metres. how high above the ground does the ladder reach?
----
You have an isosceles triangle with base = 1 meter,
each side is 2.4 meters.
----
Draw the height as a perpendicular bisector of the base.
---
Use Pythagoras to find the height.
----
slant height ^2 = height^2 + base^2
------
2.4^2 = h^2 + (1/2)^2
---
h^2 = 5.76 - 0.25
---
h^2 = 5.51
height = 2.35 meters
=========================
Cheers,
Stan H.
===========